OpenStudy (lifeisadangerousgame):

Solve for x 3x^4 + 18x^2 + 24 = 0

OpenStudy (lifeisadangerousgame):

so far I have 3(x^4 + 6x^2 + 8) = 0

OpenStudy (loser66):

3 is never =0, right? hence, the term inside the bracket =0, right?

OpenStudy (lifeisadangerousgame):

oh whoops, you're right

OpenStudy (loser66):

Now, let t = x^2, solve the quadratic in term of t

OpenStudy (lifeisadangerousgame):

t^2 + 6t + 8 so (t + 4)(t + 2)?

OpenStudy (loser66):

I think so

OpenStudy (lifeisadangerousgame):

so then really (x^2 + 4)(x^2 + 2) = 0 and solve for both?

OpenStudy (loser66):

Question: Do you have any field you have to work on? like Real/ Complex?

OpenStudy (lifeisadangerousgame):

Both at this point, I'm reviewing my third exam and seeing what I missed, we had reached complex numbers by now

OpenStudy (loser66):

ok, if it is so, you have 4 roots in complex. If not, you don't have any solution in Real

OpenStudy (lifeisadangerousgame):

I got \[i \sqrt{2}\] for both

OpenStudy (loser66):

how? for x^2+4 =0, you have 2 roots for x^2+2 =0 , you have another 2 roots They are different, right?

OpenStudy (lifeisadangerousgame):

Oh wait, I see what I did wrong

OpenStudy (lifeisadangerousgame):

x^2 + 4 = 0 -4 -4 x^2 = -4 sqrtx^2 = sqrt-4 x = \[\pm 2i\]

OpenStudy (lifeisadangerousgame):

x^2 + 2 = 0 - 2 -2 x^2 = -2 sqrtx^2 = sqrt-2 x = isqrt2?

OpenStudy (loser66):

yup for the first one, no for the second one

OpenStudy (loser66):

\(x=\pm i\sqrt{2}\)

OpenStudy (lifeisadangerousgame):

ohh okay. I wasn't sure where to put the plus/minus part, I understand that

OpenStudy (lifeisadangerousgame):

Thank you! Would you mind helping me on another one?

OpenStudy (loser66):

let see, if I can